OpenStudy (anonymous):

A line passes through the points (-10,-4) and (-1,2) . What is the slope of the line that passes through the two points?

OpenStudy (anonymous):

If by slope then you mean gradient, you can use the formula m = (y2-y1)/(x2-x1) So: m = (2 + 4)/(-1 + 10) m = 6/9 m = 2/3 Therefore the slope is 2/3

OpenStudy (anonymous):

Thanx